# Channels of Stellar-mass Black Hole Formation

### v1

@@ Abstract
On the basis of a large collection of detailed 3D core-collapse supernova simulations carried to late times, we identify four channels of stellar mass black hole formation. Our examples for Channel 1 involve the formation of lower-gap and above black holes in energetic asymmetric supernova explosions. Our Channel 2 example involves a modest supernova explosion that may leave behind a lower-gap to $\sim$10 $M_{\odot}$ black hole. The latter may not be easily distinguishable from “standard” supernovae that birth neutron stars. Our Channel 3 example experiences an aborted core-collapse explosion, more often in the context of a low-metallicity progenitor, whose residue is a black hole with a mass perhaps up to $\sim$40 $M_{\odot}$. The latter may be accompanied by a pulsational-pair instability supernova (PPISN). Channel 4 is the only quiescent or “silent” scenario for which perhaps $\sim$5 to 15 $M_{\odot}$ black holes are left. Where appropriate, we estimate 56Ni yields, explosion energies, approximate recoil speeds, and residual black hole masses. The progenitor mass density and binding energy profiles at collapse influence the outcome in a systematic way. The statistics and prevalence of these various channels depend not only on still evolving supernova theory, but on remaining issues with the theory of massive star evolution, binary interaction, wind mass loss, metallicity, and the nuclear equation of state. Importantly, we suggest, but have not proven, that the silent channel for black hole formation may not be the dominant formation modality.
@@ 1 Introduction
Stellar-mass black holes in high-mass X-ray binaries (HMXBs) and low-mass X-ray binaries (LMXBs) have been studied in the galaxy for decades (Remillard & McClintock, 2006; MacLeod & Grindlay, 2023). They have measured masses (with modest error bars) between $\sim$4 and $\sim$21 $M_{\odot}$, with the peak near $\sim$8 $M_{\odot}$ (Özel et al., 2010; Farr et al., 2011). In addition, the Gaia astrometric mission has recently found three non-accreting black holes with masses of $\sim$9 $M_{\odot}$ (BH1 and BH2, El-Badry et al., 2023a, b) and $\sim$33 $M_{\odot}$ (BH3 Gaia Collaboration et al., 2024), with many more promised. Suggestively, the lower-mass BH1 and BH2 orbit approximately solar-metallicity stars, while the more massive Gaia black hole BH3 orbits a low-metallicity ($[Fe/H]$ $\sim$ --2.5) star. Also intriguing is the fact that all three of these black holes reside in highly eccentric ($\sim$0.4, $\sim$0.5, and $\sim$0.7, respectively) orbits.
However, the recent epochal detections by the LIGO/Virgo/KAGRA (LVK) collaboration (Aasi et al., 2015; Abbott et al., 2023; Acernese et al., 2015; Akutsu et al., 2021) of gravitational waves from scores of merging black hole binaries have catapulted the study of the origin, prevalence, and mass function of such black holes to the forefront of astrophysical research (Li et al., 2024, and references therein). All these black holes are born through the collapse of the cores of massive stars, or perhaps the merger of black holes formed through such a channel, but which progenitor star for a given metallicity gives birth to stellar-mass black holes is very much a topic of ongoing research. Interestingly, the LVK collaboration has also identified in the putative lower mass gap a possible black hole in the mass interval $\sim$2.5--4.5 $M_{\odot}$ (GW230529, Abac et al., 2024). The basic questions remain: What are the detailed stellar contexts of black hole birth and the true galactic black hole mass distribution function? The number of stellar-mass black holes in the galaxy may be between $\sim$106 and $\geq$107, and this may complement an inferred 108 galactic neutron stars, but we don’t yet know.
@@
For a 40-M zero-metallicity progenitor (Woosley et al., 2010), Chan et al. (2018) and Moriya et al. (2019) witnessed a weak explosion that was accompanied by black hole birth due to significant late-time fallback of initially ejected matter. Using approximate neutrino transport and exploring initially rotating and non-rotating variants of a solar-metallicity 40-M progenitor (Sukhbold et al., 2018), Pan et al. (2021) concluded that their rotating models exploded and that the slowly-rotating and non-rotating models left black holes. The rapidly-rotating model left a neutron star. Using the same solar-metallicity 40-$M_{\odot}$ progenitor, Ott et al. (2018) found that though the core mantle initially exploded, it did so weakly and that a black hole would ultimately form without much of an external display. However, they and these other researchers calculated no model beyond $\sim$1 second after bounce, a time that has been determined to be too short to determine the asymptotic state of core collapse (Müller, 2016; Burrows et al., 2020; Bollig et al., 2021; Burrows et al., 2024a). Moreover, Walk et al. (2020), Kuroda et al. (2018), and Kuroda & Shibata (2023) observed that massive stars with ZAMS masses of 40-$M_{\odot}$ to 75-$M_{\odot}$ could launch the stalled shock, but that black hole birth rapidly followed. The short interval between explosion and black hole birth ultimately aborted neutrino heating and undermined the supernova into infall, further fattening the black hole. However, Sykes & Müller (2024a) in a series of recent 2D simulations of zero-metallicity stars from 60 to 90 $M_{\odot}$ find that they explode with canonical supernova energies, while leaving behind black holes with masses from $\sim$21 to $\sim$34 $M_{\odot}$. These could be examples of so-called “fallback” supernovae and black hole formation (see §4.2). Finally, we (Burrows et al., 2023) recently simulated in 3D the solar-metallicity 40-$M_{\odot}$ progenitor model of Sukhbold et al. (2018) using the SFHo (Steiner et al., 2013b) nuclear equation of state (EOS) and found using our state-of-the-art code Fornax that it exploded vigorously and quite aspherically.
@@ 2 Method
We employ our workhorse code Fornax to simulate the dynamical collapse and explosion phase to many seconds after bounce, starting with initial progenitor models created by KEPLER (Woosley et al., 2002; Sukhbold et al., 2016, 2018, and S. Woosley, private communication). Fornax has been exercised for the last eight years on many High-Performance-Computing (HPC) platforms and numerous publications on a wide spectrum of topics in CCSN theory have resulted (Radice et al., 2017; Burrows et al., 2018; Vartanyan et al., 2018, 2019b; Morozova et al., 2018; Burrows et al., 2019; Radice et al., 2019; Nagakura et al., 2019; Vartanyan et al., 2019a; Nagakura et al., 2020; Vartanyan & Burrows, 2020; Burrows et al., 2020; Nagakura et al., 2021; Burrows & Vartanyan, 2021; Vartanyan et al., 2022; Coleman & Burrows, 2022; Burrows et al., 2023, 2024a). A discussion of the numerical methods and microphysics employed can be found in Skinner et al. (2019), Burrows et al. (2023), and in the appendix to Vartanyan et al. (2019a). For all 3D Fornax simulations, we use the SFHo nuclear equation of state (Steiner et al., 2013a), our default spatial gridding ($1024\times 128\times 256$: $r\times\theta\times\phi$), and twelve neutrino energy groups for each of three species ($\nu_{e}$, $\bar{\nu}_{e}$, and “$\nu_{\mu}$” [$\equiv$ $\nu_{\mu}$, $\bar{\nu}_{\mu}$, $\nu_{\tau}$, $\bar{\nu}_{\tau}$]). After black hole formation (at which point the radiation module of Fornax crashes), to continue a simulation we map into either the hydrodynamic component of Fornax with a point mass and diode inner boundary condition or the FLASH code (Fryxell et al., 2000) with a similar inner boundary and point mass (Vartanyan et al., 2024). For all post-Fornax simulations, we used the Helmholtz equation of state (Timmes & Arnett, 1999). For the late-time simulation of the 23-$M_{\odot}$ model, which did not immediately form a black hole, we continued the Fornax run from 6.22 seconds after bounce with a point mass and a wind inner boundary condition (Wongwathanarat et al., 2015) at 500 kilometers (km), and then at 50 seconds after bounce mapped into FLASH and continued it for many physical days.
@@ 3 Relevant Theory of Neutrino-Driven Core-Collapse Supernova Explosions
At the end of the quasi-static evolution of stars more massive than perhaps $\sim$8 $M_{\odot}$ (which may last a few to $\sim$30 million years), their degenerate cores eventually achieve the effective Chandrasekhar mass of from $\sim$1.3 to $\sim$2.1 $M_{\odot}$ and gravitationally collapse [^3] The “Chandrasekhar” mass that becomes unstable and collapses is a function of the entropy and electron-fraction profiles achieved in the core. More massive stars evolve more quickly, thereby retaining higher core entropies that have less time to decrease due to thermal neutrino emission. This translates into more massive cores, supported by higher entropies, for more massive progenitor stars.. Within hundreds of milliseconds of instability and the onset of implosion, the central mass densities achieve values in excess of that of the atomic nucleus, the matter stiffens to near incompressibility, and the inner core bounces violently into the outer, still infalling, core mantle. At the interface, near a sonic point, a spherical shock wave is formed and launched outward. However, due predominantly to electron capture and the associated copious electron neutrino ($\nu{{}_{e}}$) energy emission behind the compressing shock wave in a “breakout neutrino burst” (and secondarily to shock photodissociation of the infalling nuclei), the shock stalls within tens of milliseconds into an accretion shock. Much theoretical work over the last decades has gone into understanding how this stalled shock is revived into explosion. The key for most models is the attendant neutrino-driven convection and turbulence behind the shock, whose primary role is to contribute turbulent stress to augment the pressure behind the shock (Burrows et al., 1995a; Janka, 2012; Burrows et al., 2020). Without this 3D ingredient most models do not explode in spherical (1D) symmetry (for exceptions, see Kitaura et al. (2006) and Wang & Burrows (2024a)). At a critical time, determined in a complicated way by the competition between neutrino heating and the augmented stress behind the shock and the ram pressure of the infalling matter (Wang et al., 2022), the structure becomes unstable (reaches a “critical point”) (Burrows & Goshy, 1993) and the shock is (re)launched. This can take between $\sim$50 and $\sim$1000 milliseconds after core bounce. The accretion of an abrupt silicon/oxygen interface can also kick a model into explosion, since upon encountering the interface the shock experiences an abrupt drop in ram pressure, while maintaining for the accretion time to the core the powering neutrino luminosity.
@@
The scenario articulated above is more or less the standard narrative and is, in broad outline, what all researchers performing detailed 3D CCSN simulations (e.g., Lentz et al., 2015; Burrows et al., 2019; Vartanyan et al., 2019a; Müller et al., 2019; Stockinger et al., 2020; Burrows et al., 2020; Bollig et al., 2021; Sandoval et al., 2021; Nakamura et al., 2022; Vartanyan et al., 2023; Burrows et al., 2024a) see and agree upon. However, there are other aspects of the explosion phenomenon that are more subtle and that are not widely appreciated, yet of central importance to the questions at hand.
@@
First, we note that many have incorrectly invoked “compactness” as an “explodability” condition (e.g., Sukhbold et al., 2016), with the suggestion that a low compactness (steep progenitor core mass density profile) is necessary for explosion by the neutrino mechanism. It isn’t, and models at both low and high compactness can explode. In fact, all our models with high compactness experience an initial (re)ignition of the stalled shock. Moreover, recent detailed simulations (Burrows et al., 2019; Vartanyan et al., 2019a; Müller et al., 2019; Stockinger et al., 2020; Burrows et al., 2020; Bollig et al., 2021; Sandoval et al., 2021; Nakamura et al., 2022; Vartanyan et al., 2023) and general theory (Bethe & Wilson, 1985; Burrows et al., 1995b; Janka, 2012; Burrows, 2013; Burrows et al., 2020) demonstrate that in the context of the neutrino heating mechanism of CCSNe a high compactness is necessary to achieve the canonical supernova energies of $\sim$1.0 Bethe (Burrows et al., 2024a; Bollig et al., 2021). High compactnesses are associated with high mass accretion rates ($\dot{M}$), that result in both high neutrino accretion luminosities and high neutrino absorption “optical depths” behind the stalled shock; the product (``$L\tau"$) is the heating power behind the shock that drives the explosion. Conversely, the lowest compactness models that easily explode444e.g., 8.8-$M_{\odot}$ (Nomoto, 1984) and $\sim$9.0 -- 9.5 $M_{\odot}$ at solar metallicity; 8.4 and 9.6 $M_{\odot}$ at very low or zero metallicity (Wang & Burrows, 2024a, and A. Heger, private communication) result in explosion energies of only $\sim$0.1 -- 0.15 Bethes (Burrows et al., 2019; Stockinger et al., 2020). The physics behind this conclusion is also discussed in Burrows & Vartanyan (2021), Wang et al. (2022), Burrows et al. (2024a), and Burrows et al. (2023). Nevertheless, and importantly, compactness is but a crude substitute for the full mass density profile of the progenitor core at collapse, but it can capture the basic trend. A few-parameter characterization, perhaps including the outer mantle binding energy and/or multiple compactnesses defined at different interior masses, would be much preferred.
@@
Third, the standing accretion shock instability (“SASI”) (Blondin et al., 2003) is a vortical-acoustic (Foglizzo, 2002) instability of the shock wave shape that can arise when the shock radius recedes or when the mass accretion rate is large. An approximate condition for this is a small “Foglizzo number” (Foglizzo et al., 2007) [^5] The Foglizzo number is proportional to the ratio of the advection time between the shock and PNS surface and the convective overturn time, itself inversely proportional to the Brunt frequency. When the shock radius is small and/or when the mass accretion rate through the shock is large, this condition can be met.. For most low-mass progenitors that explode prior to the significant recession of the shock radius, it does not appear and is completely overwhelmed by neutrino driven convection (Burrows et al., 2012; Buellet et al., 2023). In previous 2D simulations, the “sloshing” of the shock along the symmetry axis was often mistaken for the SASI, even though it was neutrino-driven convection. In 3D simulations, such axial sloshing is rarely seen. However, 1) for higher compactnesses and mass accretion rates prior to explosion; 2) at a few hundreds of milliseconds after bounce; and 3) generally for the higher mass progenitors, the mean shock radius recedes (even before explosion) and can settle from $\sim$175 kilometers to $\sim$100--125 kilometers (Burrows et al., 2020). At this point, whether or not the model explodes, the Foglizzo condition is met and a spiral SASI (Blondin & Shaw, 2007) emerges. The emergence of such a mode was also noted by Andresen et al. (2017) for a short interval around the time their 25-$M_{\odot}$ model started to explode. This has not an axial motion, but a quasi-periodic rotary motion with a period near $\sim$10 milliseconds, set approximately by the advection time between the shock and the inner PNS core (Foglizzo et al., 2007). This mode modulates the shock dipole direction, the solid-angle-integrated neutrino emissions, and the gravitational-wave signal (Andresen et al., 2017; Vartanyan et al., 2023). It is seen in all our 3D simulations that do not explode early. It is also partially responsible for the roughly monotonic-with-compactness dependence seen in the degree of ejecta asymmetry (Burrows et al., 2024a). All our 3D higher-compactness models that either explode and form neutron stars, explode and form black holes, or do not explode and later form black holes more quiescently manifest this late-time spiral SASI mode. Specifically, it is seen in all our black hole formation models in the CCSN context. Figure 2 depicts the evolution after bounce of both the mean (left) and maximum (right) shock radius with time for many of our 3D models. The cluster of higher compactness models (orange) are most delayed and clearly manifest the $\sim$100 Hz spiral SASI wobble during the launch phase.
@@
Hence, there is a race at the highest compactnesses between mass accumulation of the core and the deposition of neutrino energy [^6] This race is also dependent upon the nuclear equation of state.. For the very highest compactnesses, generally more easily achieved at low metallicities and high progenitor masses, the supernova often loses this race. The post-launch accretion rates and outer mantle binding energies are too great. Due to weaker stellar winds, for low-metallicity progenitors the corresponding final black hole masses should be higher  [^7] This expectation is consonant with the Gaia detection of the 33-$M_{\odot}$ BH3 with a low-metallicity companion (Gaia Collaboration et al., 2024)., while at solar-metallicity they should be lower. Indeed, for our highest-compactness solar-metallicity models the residual black hole masses span the range from the lower mass gap upwards. However, in all such cases the supernova is always (at least for the models we have explored) launched, and sometimes vigorously enough to result in a supernova explosion. Indeed, our 19.56- and 40-$M_{\odot}$ models have the highest asymptotic supernova energies (including the stellar mantle binding energy!) we have yet simulated.

### v2

@@ Abstract
On the basis of a large collection of detailed 3D core-collapse supernova simulations carried to late times, we identify four channels of stellar mass black hole formation. Our examples for Channel 1 can involve the formation of black holes in energetic asymmetric supernova explosions. Our Channel 2 example involves a modest supernova explosion that nevertheless leaves behind a black hole. The latter may not be easily distinguishable from “standard” supernovae that birth neutron stars. Our Channel 3 example experiences an aborted core-collapse explosion, more often in the context of a low-metallicity progenitor, whose residue is a black hole with a mass perhaps up to $\sim$40 $M_{\odot}$. The latter may be accompanied by a pulsational-pair instability supernova (PPISN). Channel 4 is the only quiescent or “silent” scenario for which perhaps $\sim$5 to 15 $M_{\odot}$ black holes are left. Where appropriate, we estimate 56Ni yields, explosion energies, approximate recoil speeds, and residual black hole masses. The progenitor mass density and binding energy profiles at collapse influence the outcome in a systematic way. We speculate that the statistics and prevalence of these various channels depend not only on still evolving supernova theory, but on remaining issues with the theory of massive star evolution, binary interaction, wind mass loss, metallicity, and the nuclear equation of state. Importantly, we suggest, but have not proven, that the silent channel for black hole formation may not be the dominant formation modality.
@@ 1 Introduction
Stellar-mass black holes in high-mass X-ray binaries (HMXBs) and low-mass X-ray binaries (LMXBs) have been studied in the galaxy for decades (Remillard & McClintock, 2006; MacLeod & Grindlay, 2023). They have measured masses (with modest error bars) between $\sim$4 and $\sim$21 $M_{\odot}$, with the peak near $\sim$8 $M_{\odot}$ (Özel et al., 2010; Farr et al., 2011). In addition, the Gaia astrometric mission has recently found three non-accreting black holes with masses of $\sim$9 $M_{\odot}$ (BH1 and BH2, El-Badry et al., 2023a, b) and $\sim$33 $M_{\odot}$ (BH3 Gaia Collaboration et al., 2024), with many more promised. These black hole masses are likely to be very close to their birth masses, since subsequent accretion in the associated wider binary context is likely to be minimal (Cappelluti et al., 2024). However, the mass added after black hole birth when in a tight binary can vary from a negligible amount to perhaps $\sim$0.5 $M_{\odot}$ (Xing, Zepei et al., 2025), but could be larger (Klencki et al., 2025). In this paper, we deal exclusively with black hole birth masses. Suggestively, the lower-mass BH1 and BH2 orbit approximately solar-metallicity stars, while the more massive Gaia black hole BH3 orbits a low-metallicity ($[Fe/H]$ $\sim$ --2.5) star. Also intriguing is the fact that all three of these black holes reside in highly eccentric ($\sim$0.4, $\sim$0.5, and $\sim$0.7, respectively) orbits.
However, the recent epochal detections by the LIGO/Virgo/KAGRA (LVK) collaboration (Aasi et al., 2015; Abbott et al., 2023; Acernese et al., 2015; Akutsu et al., 2021) of gravitational waves from scores of merging black hole binaries have catapulted the study of the origin, prevalence, and mass function of such black holes to the forefront of astrophysical research (Li et al., 2024, and references therein). All these black holes are born through the collapse of the cores of massive stars, or perhaps the merger of black holes formed through such a channel, but which progenitor star for a given metallicity gives birth to stellar-mass black holes is very much a topic of ongoing research. Interestingly, the LVK collaboration has identified in the putative lower mass gap (Shao, 2022) a possible black hole in the mass interval $\sim$2.5--4.5 $M_{\odot}$ (GW230529, Abac et al., 2024). The basic questions remain: What are the detailed stellar contexts of black hole birth and the true galactic black hole mass distribution function? The number of stellar-mass black holes in the galaxy may be as many as 108 (Kaczmarek et al., 2025), and this may complement an inferred 108 galactic neutron stars, but we don’t yet know.
@@
For a 40-M zero-metallicity progenitor (Woosley et al., 2010), Chan et al. (2018) and Moriya et al. (2019) witnessed a weak explosion that was accompanied by black hole birth due to significant late-time fallback of initially ejected matter. Using approximate neutrino transport and exploring initially rotating and non-rotating variants of a solar-metallicity 40-M progenitor (Sukhbold et al., 2018), Pan et al. (2021) concluded that their rotating models exploded and that the slowly-rotating and non-rotating models left black holes. The rapidly-rotating model left a neutron star. Using the same solar-metallicity 40-$M_{\odot}$ progenitor, Ott et al. (2018) found that though the core mantle initially exploded, it did so weakly and that a black hole would ultimately form without much of an external display. However, they and these other researchers calculated no model beyond $\sim$1 second after bounce, a time that has been determined to be too short to determine the asymptotic state of core collapse (Müller, 2016; Burrows et al., 2020; Bollig et al., 2021; Burrows et al., 2024a). Moreover, Walk et al. (2020), Kuroda et al. (2018), and Kuroda & Shibata (2023) observed that massive stars with ZAMS masses of 40-$M_{\odot}$ to 75-$M_{\odot}$ could launch the stalled shock, but that black hole birth rapidly followed. The short interval between explosion and black hole birth ultimately aborted neutrino heating and undermined the supernova into infall, further fattening the black hole  [^2] The reader is referred to Burrows et al. (2023), section 2 there, for a more in-depth discussion and critique of 3D black hole formation models of other groups, with an emphasis on the 40-$M_{\odot}$ solar-metallicity progenitor. However, Sykes & Müller (2024a) in a series of recent 2D simulations of zero-metallicity stars from 60 to 90 $M_{\odot}$ find that they explode with canonical supernova energies, while leaving behind black holes with masses from $\sim$21 to $\sim$34 $M_{\odot}$. These could be examples of so-called “fallback” supernovae and black hole formation (see §4.2). Finally, we (Burrows et al., 2023) recently simulated in 3D the solar-metallicity 40-$M_{\odot}$ progenitor model of Sukhbold et al. (2018) using the SFHo (Steiner et al., 2013b) nuclear equation of state (EOS) and found using our state-of-the-art code Fornax that it exploded vigorously and quite aspherically.
@@ 2 Method
We employ our workhorse code Fornax to simulate the dynamical collapse and explosion phase to many seconds after bounce, starting with initial progenitor models created by KEPLER (Woosley et al., 2002; Sukhbold et al., 2016, 2018, and S. Woosley, private communication). Fornax has been exercised for the last eight years on many High-Performance-Computing (HPC) platforms and numerous publications on a wide spectrum of topics in CCSN theory have resulted (Radice et al., 2017; Burrows et al., 2018; Vartanyan et al., 2018, 2019b; Morozova et al., 2018; Burrows et al., 2019; Radice et al., 2019; Nagakura et al., 2019; Vartanyan et al., 2019a; Nagakura et al., 2020; Vartanyan & Burrows, 2020; Burrows et al., 2020; Nagakura et al., 2021; Burrows & Vartanyan, 2021; Vartanyan et al., 2022; Coleman & Burrows, 2022; Burrows et al., 2023, 2024a). A discussion of the numerical methods and microphysics employed can be found in Skinner et al. (2019), Burrows et al. (2023), and in the appendix to Vartanyan et al. (2019a). For all 3D Fornax simulations, we use the SFHo nuclear equation of state (Steiner et al., 2013a), our default spatial gridding ($1024\times 128\times 256$: $r\times\theta\times\phi$), and twelve neutrino energy groups for each of three species ($\nu_{e}$, $\bar{\nu}_{e}$, and “$\nu_{\mu}$” [$\equiv$ $\nu_{\mu}$, $\bar{\nu}_{\mu}$, $\nu_{\tau}$, $\bar{\nu}_{\tau}$]). After black hole formation (when the lapse iteration does not converge, and the code crashes due to the inability to find a stable solution consistent with general relativity (see below), to continue a simulation we map into either the hydrodynamic component of Fornax with a point mass and diode inner boundary condition or the FLASH code (Fryxell et al., 2000) with a similar inner boundary and point mass (Vartanyan et al., 2024). For all post-Fornax simulations, we used the Helmholtz equation of state (Timmes & Arnett, 1999). For the late-time simulation of the 23-$M_{\odot}$ model, which did not immediately form a black hole, we continued the Fornax run from 6.22 seconds after bounce with a point mass and a wind inner boundary condition (Wongwathanarat et al., 2015) at 500 kilometers (km), and then at 50 seconds after bounce mapped into FLASH and continued it for many physical days. The wind was constructed from the flow and its properties actually witnessed in the corresponding Fornax simulation and its time behavior was extrapolated using a power law fit (see, Wang & Burrows, 2024a). The juncture was exterior to the sonic point of the flow. In this way there is a smooth hydrodynamic transition at the code hand-off, with no artificial shocks or discontinuities.
Fornax incorporates approximate general relativity, with gravitational redshifts applied to the radiation fields (Vartanyan et al., 2019a). The monopole term of the gravitational field is the same as in the relativistic TOV equation, with the neutrino energy density and pressure, as well as the kinetic term, included as described in Marek et al. (2006), Case A. In this way, if the core were perfectly spherical the effect of general relativity on the quasi-hydrostatic inner structure would be accurate. The lapse function is derived from this function and directly yields the redshifts. An iterative calculation is required to obtain the gravitational masses and this iteration does not converge when the core is unstable to gravitational collapse. This is also the approach used in Walk et al. (2020). Plots of the lapse at an early and terminal phase can be found in Burrows et al. (2023).
@@ 3 Relevant Theory of Neutrino-Driven Core-Collapse Supernova Explosions
At the end of the quasi-static evolution of stars more massive than perhaps $\sim$8 $M_{\odot}$ (which may last a few to $\sim$30 million years), their degenerate cores eventually achieve the effective Chandrasekhar mass of from $\sim$1.3 to $\sim$2.1 $M_{\odot}$ and gravitationally collapse [^4] The “Chandrasekhar” mass that becomes unstable and collapses is a function of the entropy and electron-fraction profiles achieved in the core. More massive stars evolve more quickly, thereby retaining higher core entropies that have less time to decrease due to thermal neutrino emission. This translates into more massive cores, supported by higher entropies, for more massive progenitor stars.. Within hundreds of milliseconds of instability and the onset of implosion, the central mass densities achieve values in excess of that of the atomic nucleus, the matter stiffens to near incompressibility, and the inner core bounces violently into the outer, still infalling, core mantle. At the interface, near a sonic point, a spherical shock wave is formed and launched outward. However, due predominantly to electron capture and the associated copious electron neutrino ($\nu{{}_{e}}$) energy emission behind the compressing shock wave in a “breakout neutrino burst” (and secondarily to shock photodissociation of the infalling nuclei), the shock stalls within tens of milliseconds into an accretion shock. Much theoretical work over the last decades has gone into understanding how this stalled shock is revived into explosion. The key for most models is the attendant neutrino-driven convection and turbulence behind the shock, whose primary role is to contribute turbulent stress to augment the pressure behind the shock (Burrows et al., 1995a; Janka, 2012; Burrows et al., 2020). Without this 3D ingredient most models do not explode in spherical (1D) symmetry (for exceptions, see Kitaura et al. (2006) and Wang & Burrows (2024b)). At a critical time, determined in a complicated way by the competition between neutrino heating and the augmented stress behind the shock and the ram pressure of the infalling matter (Wang et al., 2022), the structure becomes unstable (reaches a “critical point”) (Burrows & Goshy, 1993) and the shock is (re)launched. This can take between $\sim$50 and $\sim$1000 milliseconds after core bounce. The accretion of an abrupt silicon/oxygen interface can also kick a model into explosion, since upon encountering the interface the shock experiences an abrupt drop in ram pressure, while maintaining for the accretion time to the core the powering neutrino luminosity.
@@
The scenario articulated above is more or less the standard narrative and is, in broad outline, what all researchers performing detailed 3D CCSN simulations (Lentz et al., 2015; Burrows et al., 2019; Vartanyan et al., 2019a; Müller et al., 2019; Stockinger et al., 2020; Burrows et al., 2020; Bollig et al., 2021; Sandoval et al., 2021; Nakamura et al., 2022; Vartanyan et al., 2023; Burrows et al., 2024a) see and agree upon. However, there are other aspects of the explosion phenomenon that are more subtle and that are not widely appreciated, yet of central importance to the questions at hand.
@@
First, we note that many have incorrectly invoked “compactness” as an “explodability” condition (e.g., Sukhbold et al., 2016), with the suggestion that a low compactness (steep progenitor core mass density profile) is necessary for explosion by the neutrino mechanism. It isn’t, and models at both low and high compactness can explode. In fact, all our models with high compactness experience an initial (re)ignition of the stalled shock. Moreover, recent detailed simulations (Burrows et al., 2019; Vartanyan et al., 2019a; Müller et al., 2019; Stockinger et al., 2020; Burrows et al., 2020; Bollig et al., 2021; Sandoval et al., 2021; Nakamura et al., 2022; Vartanyan et al., 2023) and general theory (Bethe & Wilson, 1985; Burrows et al., 1995b; Janka, 2012; Burrows, 2013; Burrows et al., 2020) demonstrate that in the context of the neutrino heating mechanism of CCSNe a high compactness is necessary to achieve the canonical supernova energies of $\sim$1.0 Bethe (Burrows et al., 2024a; Bollig et al., 2021). High compactnesses are associated with high mass accretion rates ($\dot{M}$), that result in both high neutrino accretion luminosities and high neutrino absorption “optical depths” behind the stalled shock; the product (``$L\tau"$) is the heating power behind the shock that drives the explosion. Conversely, the lowest compactness models that easily explode555e.g., 8.8-$M_{\odot}$ (Nomoto, 1984) and $\sim$9.0 -- 9.5 $M_{\odot}$ at solar metallicity; 8.4 and 9.6 $M_{\odot}$ at very low or zero metallicity (Wang & Burrows, 2024b, and A. Heger, private communication) result in explosion energies of only $\sim$0.1 -- 0.15 Bethes (Burrows et al., 2019; Stockinger et al., 2020). The physics behind this conclusion is also discussed in Burrows & Vartanyan (2021), Wang et al. (2022), Burrows et al. (2024a), and Burrows et al. (2023). Nevertheless, and importantly, compactness is but a crude substitute for the full mass density profile of the progenitor core at collapse, but it can capture the basic trend. A few-parameter characterization, perhaps including the outer mantle binding energy and/or multiple compactnesses defined at different interior masses, would be much preferred.
@@
Figure 3: This figure depicts the three dipole components of the spherical harmonic decomposition of the shock radius as a function of time for the 20 $M_{\odot}$ model. The $\sim$100 Hz spiral-SASI mode becomes significant at around 0.4 seconds post bounce and lasts for about 0.3 seconds until the model explodes at about 0.7 seconds post bounce (See Figure 2).
@@
Third, the standing accretion shock instability (“SASI”) (Blondin et al., 2003) is a vortical-acoustic (Foglizzo, 2002) instability of the shock wave shape that can arise when the shock radius recedes or when the mass accretion rate is large. An approximate condition for this is a small “Foglizzo number” (Foglizzo et al., 2007) [^6] The Foglizzo number is proportional to the ratio of the advection time between the shock and PNS surface and the convective overturn time, itself inversely proportional to the Brunt frequency. When the shock radius is small and/or when the mass accretion rate through the shock is large, this condition can be met.. For most low-mass progenitors that explode prior to the significant recession of the shock radius, it does not appear and is completely overwhelmed by neutrino driven convection (Burrows et al., 2012; Buellet et al., 2023). In previous 2D simulations, the “sloshing” of the shock along the symmetry axis was often mistaken for the SASI, even though it was neutrino-driven convection. In 3D simulations, such axial sloshing is rarely seen. However, 1) for higher compactnesses and mass accretion rates prior to explosion; 2) at a few hundreds of milliseconds after bounce; and 3) generally for the higher mass progenitors, the mean shock radius recedes (even before explosion) and can settle from $\sim$175 kilometers to $\sim$100--125 kilometers (Burrows et al., 2020). At this point, whether or not the model explodes, the Foglizzo condition is met and a spiral SASI (Blondin & Shaw, 2007) emerges. The emergence of such a mode was also noted by Andresen et al. (2017) for a short interval around the time their 25-$M_{\odot}$ model started to explode. This has not an axial motion, but a quasi-periodic rotary motion with a period near $\sim$10 milliseconds, set approximately by the advection time between the shock and the inner PNS core (Foglizzo et al., 2007). This mode modulates the shock dipole direction, the solid-angle-integrated neutrino emissions, and the gravitational-wave signal (Andresen et al., 2017; Vartanyan et al., 2023). It is seen in all our 3D simulations that do not explode early. It is also partially responsible for the roughly monotonic-with-compactness dependence seen in the degree of ejecta asymmetry (Burrows et al., 2024a). All our 3D higher-compactness models that either explode and form neutron stars, explode and form black holes, or do not explode and later form black holes more quiescently manifest this late-time spiral SASI mode. Specifically, it is seen in all our black hole formation models in the CCSN context. Figure 2 depicts the evolution after bounce of both the mean (left) and maximum (right) shock radius with time for many of our 3D models. The cluster of higher compactness models (orange) are most delayed and clearly manifest the $\sim$100 Hz spiral SASI wobble during the launch phase. This spiral SASI behavior is more clearly seen in Figure 3, which depicts the temporal variation of the dipole components of the shock surface for the 20 $M_{\odot}$ model.
@@
Hence, there is a race at the highest compactnesses between mass accumulation of the core and the deposition of neutrino energy [^7] This race is also dependent upon the nuclear equation of state.. For the very highest compactnesses, generally more easily achieved at low metallicities and high progenitor masses, the supernova often loses this race. The post-launch accretion rates and outer mantle binding energies are too great. Due to weaker stellar winds, for low-metallicity progenitors the corresponding final black hole masses should be higher  [^8] This expectation is consonant with the Gaia detection of the 33-$M_{\odot}$ BH3 with a low-metallicity companion (Gaia Collaboration et al., 2024)., while at solar-metallicity they should be lower. Indeed, for our highest-compactness solar-metallicity models the residual black hole masses span a range. However, in all such cases the supernova is always (at least for the models we have explored) launched, and sometimes vigorously enough to result in a supernova explosion. Indeed, our 19.56- and 40-$M_{\odot}$ models have the highest asymptotic supernova energies (including the stellar mantle binding energy!) we have yet simulated.